A company does not plan to pay dividends for the next three years. After that, they will make a fixed dividend payment of $5 per year for 5 years. After that, they will increase the dividend by 1% per year indefinitely. The required return on the stock is 14%. What is the value of the stock today? $25.20 What is the capital gains yield this year? 19.8% What is the dividend yield this year? 14% What is the value of the stock in 8 years? $71.90 What is the capital gains yield in 10 years? 1% What is the dividend yield in 13 years? 14% What is the value of the stock in 3 years? $26.49 What is the dividend yield in 3 years? 13.00% What is the capital gains yield in 3 years? 0.61%

correct ans:

a) 25.20, b) 14%, c) 0%, d) 38.85, e) 1%, f) 13%, g) 37.34,h) 13.39, I) 0.61%
Answer #1

1.
Price today=5/1.14^4+5/1.14^5+5/1.14^6+5/1.14^7+5/1.14^8+(5*1.01/(14%-1%))/1.14^8=25.20
3.
Dividend yield=0%
2.
Capital gains yield=required return-dividend yield=14%-0%=14%
4.
Value of the stock in 8 years=(5*1.01/(14%-1%))=38.85
5.
Price in 10 years=5*1.01^3/(14%-1%)
Dividend yield in 13 years=5*1.01^3/(5*1.01^3/(14%-1%))=13%
Capital gains yield in 10 years=14%-13%=1%
6.
Price in 13 years=5*1.01^6/(14%-1%)
Dividend yield in 13 years=5*1.01^6/(5*1.01^6/(14%-1%))=13%
7.
Value in 3 years=5/1.14^1+5/1.14^2+5/1.14^3+5/1.14^4+5/1.14^5+(5*1.01/(14%-1%))/1.14^5=37.34
8.
Dividend yield in 3 years=5/37.34=13.39%
9.
Capital gains yield in 3 years=14%-13.39%=0.61%